Systems Staff Meeting Notes 4/26/99

Systems Staff Meeting Notes
April 26, 1999

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Suzy Shaw, Ying Tang
Announcements
  1. UF has apparently hired a new Chief Information Officer, Mr. Earl Robbins. There has been talk of this position for a while - it is well known that it is difficult for Lombardi to deal with different people in charge of different realms of our techno fiefdom. What degree of coordination & authority this position will have is unknown, but it is very high ranking - comparable to a Senior Vice President. There is concern that all UF DP titles could be consolidated under one control. There are many work environments that are set up that way, but whether or not that is our fate is unknown. Earl has most recently been the chief technoid at the Registrar's Office and is the designer of the ISIS system.
  2. Systems Minutes visibility - Bill received an email from a woman in upstate New York who had been reading our minutes (accessed as the result of a search) and had a question about a type of software that we had been discussing at one of our meetings!
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 40 Cows should be in soon. They will be housed in our vaults. We will transfer our surveyed equipment to room 419 LAC to make space in the vaults.
  2. An additional 60 Cows will be ordered to replace the older 75 & 100 models. This purchase is library funded.
  3. 35-50 network switches will be ordered. These will replace all of our hubs and put everyone on switch port access. This will allow for much more bandwidth per work station. The collision problem in Preservation disappeared when a switch was installed. We'll be trading in the hubs, getting back more on the hubs by a factor of two then what we paid for them (rebated at close to half the purchase price).
  4. A work order has gone out to rekey the Systems front door. When this happens, the same key that works for all of our other doors will work there and the key that currently works for the front door won't. The old key will need to be put in an envelope with your name on it and returned to Bill.
Networking
  1. The listserv mail problem we experienced last week seems to have been a hang up at the University of Georgia.
  2. A patron came in to the office today, reporting that someone had used his email & was impersonating him. He wanted us to check our records to determine who may be the culprit. Unfortunately we are not able to make that determination, but Bill has turned the case over to the Threat Response Group.
  3. Proxy Server - has been operating fine since the initial glitch. The RAM strain is a bit high, but the filtering is working extremely well. Bill recently blocked access to the directory of a free email address site, as well as Alta Vista & Yahoo, because someone figured a way around the recent filter at those sites. It is a continual cat & mouse game.
  4. Chernobyl Virus Threat - this is an old virus - since we've done the full virus scan upgrade with the threat of Melissa, we should be okay. We have no reports of it in the Libraries. It is, however, a nasty virus that attacks the flash bios and it should be avoided.
  5. Wiring - there is still a hold up on the racks. The fiber has been terminated in MSL. East is complete, AFA not yet started. Webmaster Suite is finished & the phone work has been done. It has not been furnished yet. Punch downs in the closet are done, the suite will be on our phone roll over.
  6. Hang is doing some wiring to partitions in Access Services. We'll need to run wires for them. We also need to do a little retro work in the back of Access Services, raising the wires for them.
Software Projects Status
  1. ATC standalone catalog - Winston is having a keywords problem with it. He determined that it's not a linear system, rather exponential. The process needs to be able to be reproduced. Winston will try some other avenues. The fact is, the ATC is functional without keywords. For short network outages, it really would not even be necessary. We may just pass on the keyword function.
  2. LASSIE - have been trying to load old information to it. There has been no fund information entered on past orders. This will require a different mechanism. Currently using dummy pay sources.
  3. Winston & Brian are doing some requirements research for Betty Mitchell. She is interested in keeping track of orders & generating reports from them. Betty also mentioned to Brian that they are getting a new VISA card & that's Scott's access on it is in question.
  4. Requirement research is also underway for LAC - they are interested in a way to monitor patrons coming in & out of the stacks. They would like to use the barcode reader. It is unclear if this is ethically possible.
Calendar
  1. Everett out today, April 26. Bobbie is also out (sick).
  2. Ashley out Wednesday, April 28.
  3. Rich out April 30-May 3.
  4. Brian out May 13 & 14.
  5. Rich out May 14.
  6. Brian out May 28-June 1.
  7. Bobbie out July 12-July 16.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu